OpenStudy (anonymous):

if f(x)= (2x+1)^4 than what is the fourth derivative?

OpenStudy (loser66):

take the first then the second, then the third then the fourth NO shortcut. kha kha

OpenStudy (anonymous):

can you explain how to do the first one, I think i am messing up the chain rule part

OpenStudy (loser66):

(2x+1)^4 the first derivative = 4((2x+1)^3* 2

OpenStudy (loser66):

=8(2x+1)^3

OpenStudy (anonymous):

so the second derivative would then be 48(2x+1)^2 ?

OpenStudy (loser66):

yup

OpenStudy (anonymous):

and nothing on the inside changes?

OpenStudy (loser66):

yup, fortunately, the inside term is just degree of 1. if its degree is 5, you are gonna go crazy khakhakha

OpenStudy (anonymous):

so when you go from the 3rd derivative to the 4th derivative: f^iii(x)= 192(2x+1) to f^iv(x)=384(2x+1)... is that right?

OpenStudy (loser66):

nope, the 3rd is ok, but the 4th is not

OpenStudy (anonymous):

so what happens with the parenthesis?

OpenStudy (loser66):

Actually, it is I am sorry, I must go back and explain you why \[(2x+1)^4)'= 4(2x+1)^3*\color{red}{(2x+1)'}\] and \(\color{red}{(2x+1)'}=2\) that's why the first derivative =\( 8 (2x+1)^3\)

OpenStudy (loser66):

now, the second derivative \[(8(2x+1)^3)'=8*3(2x+1)^2*\color{red}{(2x+1)'}\] again, it is =2, so that the second derivative = \(48(2x+1)^2\)

OpenStudy (anonymous):

so then is the 4th derivative just a number and you just use the derivative of (2x+1), or 2?

OpenStudy (loser66):

the third one = 192(2x+1)= 394x + 394 so the fourth one is just 394, no more x

OpenStudy (anonymous):

ah, you factor it out and then take the derivative. I understand now. Thank you

OpenStudy (loser66):

ok